Mr. Ben bought some copies of book at GHC 5.00 each. 12 of the books got missing he then sold the rest at GHC 7.00 each and made a profit of GHC 150.00. How many books did he buys?

Answer by ikleyn(53763)

.
Mr. Ben bought some copies of book at GHC 5.00 each.
12 of the books got missing he then sold the rest at GHC 7.00 each
and made a profit of GHC 150.00. How many books did he buys?
~~~~~~~~~~~~~~~~

Let n be the number of bought copies;  then the number of sold copies was (n-12).


The profit equation is


    7*(n-12) - 5n = 150.


Simplify and solve.


    7n - 84 - 5n = 150

        2n       = 150 + 84

        2n       = 234

         n       = 234/2 = 117.      ANSWER

Solved.